Determine the amplitude, period, and phase shift of the function y=1/2sin (x+pi)
and sketch one complete period of the graph of the function..
My Answer: Amplitude=1/2 Period=2pi Phase shift=pi/3
I'm having trouble understanding how to do the rest...thanks for your help in advance..

Answer by stanbon(75887)

Determine the amplitude, period, and phase shift of the function
y=(1/2)sin (x+pi)
and sketch one complete period of the graph of the function..
My Answer: Amplitude=1/2 Period=2pi Phase shift=pi/3
------------------
a=(1/2) ; b=1 ; c=pi
Amp = abs(a) = 1/2
Period =2pi/b = 2pi
Phase = -c/b = -pi/1 = -pi
-------------------------------
Comment:
For y=sinx the sin curve period starts at theta=0 and finishes at theta=2pi
---------
But you have a phase shift of -pi so the period starts at theta = pi and
ends at 3pi.
Once you have sketched and labeled that period you can repeat it to the
right and to the left to sketch as much of the function as you wish.
--------------
Cheers,
Stan H.